A software designer is mapping the streets for a new racing game. All of the streets are depicted as either perpendicular or parallel lines. The equation of the lane passing through A and B is -7x + 3y = -21.5. What is the equation of the central street PQ?




1
Expert's answer
2021-06-16T12:23:59-0400

The equation of the line passing through A and B is

"-7x+3y=-21.5"

This could be written as

"y=\frac{7}{3}x-\frac{21.5}{3}"

By comparing with the standard form "y=mx+c"

The slope "m=\frac{7}{3}"

Central street PQ will be perpendicular to the lane passing through A and B

Product of slope of perpendicular lines will be equal to -1

"m_1.m_2=-1"

slope of perpendicular line

"=\frac{-1}{slope\space parallel\space line}=\frac{-1}{\frac{7}{3}}=\frac{-3}{7}"

This is the slope of the line.

Thus the equation of line will be "y=mx+c"

"y=\frac{-3}{7}x+c"

"7y=-3x+7c\\7y+3x=7c"


dividing by 2

"3.5y+1.5x=3.5c"


To find c with this information is impossible

To check for the line with slope as "\frac{-3}{7}" OP in a similar figure is found




The line PQ passes through the point(7,6) in the figure above hence the line could be found out using point slope form of line.

"y-y_1=m(x-x_1)"

"slope=\frac{-3}{7} (7,6)"

"y-6=\frac{-3}{7}(x-7)\implies7(y-6)=-3(x-7)\\7y-42=-3x+21\implies7y+3x=21+42"

"7y+3x=63" is the equation of central lane

Dividing by 2

"\implies3.5y-1.5x=-31.5" is also the equation.